OpenStudy (anonymous):

how do you simplify this? x^2+x-30/4x^2-100

OpenStudy (anonymous):

x^2+6x-5x-3o/4(x^2-5^2)=(x-5)(x+6)/4(x+5)(x-5)=(x+6)/4(x+5)